OpenStudy (ttop0816):

Express the Cartesian coordinates (-2,2√3) as polar coordinates. (-4, 30°) (4, -60°) (4, 120°) (4, 150°) <--- is this right??

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

4,120

OpenStudy (anonymous):

how did u find 4

OpenStudy (sidsiddhartha):

\[take ~~x=rcos \theta~~and \\y=r \sin \theta\] now square them and add

OpenStudy (sidsiddhartha):

u'll get \[r^2=x^2+y^2\\r=\sqrt{x^2+y^2}=\sqrt{16}=4\]

OpenStudy (anonymous):

then I did sine